OpenStudy (adunb8):

HELP! linear algebra & differential question: looking for exact equation ex: (y^-3-y^-2sinx)y'+y^-1cosx = 0 my final solution should be y = ?

OpenStudy (unklerhaukus):

\[\left(y^{-3}-y^{-2}\sin (x)\right)y'+y^{-1}\cos (x) = 0\]

OpenStudy (unklerhaukus):

is this your equation?

OpenStudy (adunb8):

yes

OpenStudy (lgbasallote):

and you're looking for EXACT??

OpenStudy (adunb8):

yes!

OpenStudy (kainui):

I'm reading this: http://tutorial.math.lamar.edu/Classes/DE/Exact.aspx I'll race you to see who can solve it the fastest!

OpenStudy (anonymous):

multiply through by y and it's pretty easy

OpenStudy (adunb8):

well obviously if you know it then its easy but i dont so its hard and your reply is very vague...

OpenStudy (anonymous):

multiply through by y you'll have something like f(x,y) *dy/dx + f(x) =0

OpenStudy (anonymous):

you use the exact solution method that he linked

OpenStudy (kainui):

So basically by reading this I found out that you have it in a form that's \[\Psi _{x} + \Psi _{y}\frac{ dy }{ dx }=0\] So what you have is the derivative of a function. Now you need to take the derivative of each derivative with respect to the other function to see if they're equal so you know they're an exact equation.

OpenStudy (anonymous):

f(x,y) is the partial of some function (the solution) with respect to y so integrate to find F(x,y) f(x,y) = d ( F(x,y) ) / dy

OpenStudy (kainui):

You'll find that formula I wrote on your own (I did) by using implicit differentiation. It's multivariable calculus but it's not too bad, I can help you out but keep in mind I'm still figuring this out right now lol.

OpenStudy (adunb8):

this is what i did im trying to solve what y will equal ..

OpenStudy (adunb8):

any help? i just need to solve for y but im not getting what my teacher put on his answer sheet.

OpenStudy (anonymous):

I think you're done...

OpenStudy (anonymous):

what was your question? you have it all worked out there on the page...

OpenStudy (kainui):

I know how to find y, you multiply your final answer by y on both sides until you can do the quadratic equations.

OpenStudy (kainui):

lol nvm sinx as a coefficient doesn't work, what am I thinking? I need sleep.

OpenStudy (anonymous):

what did your teacher get?

OpenStudy (kainui):

I don't know how to solve for y in this case, kinda funky to me.

OpenStudy (anonymous):

multiply by y simplify

OpenStudy (kainui):

Try it out, it doesn't go so easily I'm afraid.

OpenStudy (anonymous):

I meant in the beginning, before you do the 'exact soln.' method.... but he already did the longer version... and now he's done, but evidently his teacher got a different answer...

OpenStudy (anonymous):

I would like to know what the teacher's answer is, because his work looks correct.

OpenStudy (kainui):

\[\frac{ \sin(x) }{ y } = \frac{ 1 }{ 2y^2 }+C\]

OpenStudy (anonymous):

so he forgot the 2 when integrating to find g(y)

OpenStudy (adunb8):

basically the aswer should be \[y = [sinx \pm(\sin^2x+c)^1/2]^-1\]

OpenStudy (kainui):

and the negative sign.

OpenStudy (anonymous):

wut

OpenStudy (kainui):

Is that allowed adunb8? I guess when I said sin(x) can't be a coefficient I was wrong.

OpenStudy (anonymous):

does that even say.

OpenStudy (kainui):

It's the quadratic formula with y in there. Makes sense to me, I stopped myself because it didn't make sense when simplifying it but really we just needed an equation for y. Good job.

OpenStudy (adunb8):

iono how he got this!!! he a math genius but wont tell us how he got it!

OpenStudy (anonymous):

looks like he just used the quadratic formula after multiplying through by y^2

OpenStudy (anonymous):

Yeah, what @Kainui said...

OpenStudy (adunb8):

oh i would use quadratic formula to solve this one?

OpenStudy (kainui):

\[C+y^{-1}sinx-\frac{ 1 }{ 2 }y^{-2}=0\] now multiply each side by y^2, totally allowed in algebra. You end up with a=C, b=sinx, and c=-1/2 which are the weirdest coefficients I've ever plugged into the quadratic formula lol.

OpenStudy (anonymous):

multiply by y^2 and plug the coefficients on y^2 y etc into the quad. formula

OpenStudy (adunb8):

omg how am i suppose to know this for a test =(

OpenStudy (adunb8):

@Kainui where did you get the -1/2 from?

OpenStudy (anonymous):

@adunb8 fyi, if you have the problem done and you just have a procedural question, you can say that at the beginning and save yourself and everyone else a lot of time.

OpenStudy (anonymous):

@adunb8 go back to your integration to find g(y) and check it...

OpenStudy (adunb8):

oh.. sorry... well i still needed the refreshment on the exact its still confusing even still..

OpenStudy (adunb8):

it was still very helpful for me to look at all the steps.

OpenStudy (kainui):

You're in diffeq and don't know the quadratic formula yet...? eiiieee.... also the -1/2 came from integrating y^-2.

OpenStudy (anonymous):

you'll see the coeff. is supposed to be -1/2

OpenStudy (adunb8):

i mean i know how to do it but its confusing when he puts it like this and i gotta solve it from scratch..

OpenStudy (kainui):

I literally learned how to solve this problem from that link I showed you after I posted it and found it on google. I've only taken up to calculus II, lol.

OpenStudy (kainui):

That website is awesome lol.

OpenStudy (anonymous):

It is a great great site, tons of stuff on there.

OpenStudy (kainui):

Earlier when I said I was racing, I was literally racing you lol. XD Well any more questions or am I off to bed?

OpenStudy (adunb8):

no thank you so much for your help =) geez im a buzzkill

OpenStudy (anonymous):

s'ok, we love you <3

OpenStudy (adunb8):

=)

OpenStudy (anonymous):

:)